OpenStudy (abhisar):

Two point charges -20 C and 40 C are placed r distance apart, then potential due to these charges will be zero in ?

OpenStudy (abhisar):



OpenStudy (abhisar):

@Vincent-Lyon.Fr

OpenStudy (abhisar):

How to mathematically show that the potential will be zero in region A and B ?

OpenStudy (vincent-lyon.fr):

You need to be twice as far away from the 40 C charge as from the -20 C charge. It is not possible in region C where the first former distance is always smaller than the latter. Actually, in the plane, the solution will be given by two hyperbolas.

OpenStudy (abhisar):

Actually i want to know how u deduced the above conclusion ?

OpenStudy (abhisar):

This one You need to be twice as far away from the 40 C charge as from the -20 C charge.

OpenStudy (abhisar):

Like this, \(\large \sf \frac{k\times40}{r}=\frac{k\times20}{d}\\ --> r=2d\) ??

OpenStudy (abhisar):

Oh so what you wr trying to say that null point will always be closer to the smaller charge i.e -20C. In the region C any point will be closer to the 40C charge and farther from the -20C charge..?

OpenStudy (vincent-lyon.fr):

Yes! \(V=\large k\frac{q_1}{r_1}+k \frac{q_2}{r_2}\)

OpenStudy (abhisar):

yeah..yeah...i got it now ! Thank you very much sir ! \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)